Calculate the current, I, through the batteries for:
- a 2-bulb parallel circuit
- How does the total circuit conpare to the current through each bulb?
-a 3-bulb parallel circuit?

GIVEN THAT: The bulbs used are rated at 0.30Amps, 2.5 Volts, and 8.33 ohms.

Would the answer be .3Amps for any of the parallel circuits?
Answer #1

In parallel circuit, currents are not same usually. However, currents will be same if resistances of bulbs and voltage across them is same.

from ohm's law

V = IR

if V = 2.5 V and R = 8.33 then

I = 2.5 / 8.33

I = 0.3 A

so,

both bulbs in parallel will have same current (0.3 A ) flowing through them. However, the current flowing from the battery will be 0.3 + 0.3 = 0.6 A

In same way, for 3 bulbs in parallel, current flowing from battery will be 0.3 + 0.3 + 0.3 = 0.9 A .